How long does it take for 900 to accumulate to 1000 under an interest rate of 4% per annum?


1
Expert's answer
2022-03-10T17:58:56-0500

"A=P(1+rt)"


"1000=900(1+0.04\u00d7t)"


"\frac{1000}{900}=1+0.04t"


"\frac{10}{9}-1=0.04t"


"\frac{1}{9}=0.04t"


"t=\frac{1}{9}÷0.04=2.78 years"